OpenStudy (anonymous):

At 10:00 a.m. two planes leave an airport. If the northbound plane flies at 280 mph and the southbound at 320 mph, at what time will they be 1,000 miles apart? Which equation can be used to solve for t, the time it takes the planes to be 1,000 miles apart?

OpenStudy (anonymous):

At 10:00 a.m. two planes leave an airport. If the northbound plane flies at 280 mph and the southbound at 320 mph, at what time will they be 1,000 miles apart? Which equation can be used to solve for t, the time it takes the planes to be 1,000 miles apart? ----------------- Their speeds are added. 280 + 320 = 600 mi/hr ---------- d = r*t t = d/r

OpenStudy (anonymous):

does that help at all?